OpenStudy (rock_mit182):

Jane drove 260 miles using 12 gallons of gas. At this rate, how many gallons of gas would she need to drive 286 miles?

OpenStudy (rock_mit182):

i cant fail on this one, i mean is the last question

OpenStudy (anonymous):

\[\frac{12}{260}=\frac{x}{286}\iff x=\frac{286\times 12}{260}\]

OpenStudy (anonymous):